OpenStudy (wade123):

Eliminate the parameter. x = 3 cos t, y = 3 sin t im so confused, can someone please explain and tell me the steps

OpenStudy (wade123):

can you show me the work? i dont really understand what you mean. I have the answer, I just dont understand how to solve it @mayankdevnani

OpenStudy (wade123):

yes

OpenStudy (wade123):

@mayankdevnani

OpenStudy (wade123):

did you write something?.. @mayankdevnani

OpenStudy (mayankdevnani):

@wade123 just divide both equationss because to cancel out 3:- we get, \[\huge \sf \frac{x}{y}=\frac{3cost}{3sint}=tant\] finally we get, \[\huge \sf x=y~tant\]

OpenStudy (mayankdevnani):

understood? @wade123

OpenStudy (wade123):

you said something before about equaling it to t.. @mayankdevnani

OpenStudy (mayankdevnani):

noo...just forgot my comment and just concentrate in my new comment and if you have any problem regarding my new comment,please reply

OpenStudy (wade123):

what is tant... @mayankdevnani

OpenStudy (wade123):

@ganeshie8 please help!

OpenStudy (mayankdevnani):

\[\huge \sf tant=\frac{sint}{cost}\]

OpenStudy (mayankdevnani):

its wrong.......its \[\huge \sf \cot t\]

ganeshie8 (ganeshie8):

this is tricky, `t` is still there, try this may be x = 3 cos t, y = 3 sin t x^2 = 9 cos^2t y^2 = 9 sin^2t -------------------- x^2 + y^2 = 9(sin^2t + cos^2t)

ganeshie8 (ganeshie8):

recall the identity : sin^2t + cos^2t=1 x^2 + y^2 = 9(1) x^2 + y^2 = 9 this is just a circle !

OpenStudy (mayankdevnani):

@ganeshie8 what about my answer??

ganeshie8 (ganeshie8):

see if that makes more or less sense..

ganeshie8 (ganeshie8):

which answer ?

OpenStudy (mayankdevnani):

\[\huge \sf x=y \cot t\]

ganeshie8 (ganeshie8):

`t` still exists right ?

ganeshie8 (ganeshie8):

in the final equation, `t` should not be present...

OpenStudy (mayankdevnani):

ohhh